OpenStudy (narissa):

what is the value of x;12-3/5x-1/5x=-16

OpenStudy (anonymous):

This one is tricky

OpenStudy (narissa):

yes

OpenStudy (anonymous):

you're using / as divide right?

OpenStudy (narissa):

i have 3 more after this one after im finished can u check the first four that i did myself

OpenStudy (narissa):

its fraction

OpenStudy (anonymous):

sure. Are you in Algebra 1?

OpenStudy (narissa):

three fifths and one third

OpenStudy (narissa):

ooops one fifth

OpenStudy (narissa):

yes

OpenStudy (anonymous):

is this A1?

OpenStudy (anonymous):

I'm not that great with fractions honestly but I can try.

OpenStudy (narissa):

lol the book says math plus second edition

OpenStudy (anonymous):

12-3/5x-1/5x=-16 -3/5x-1/5x=-28 -4/5x=-28

OpenStudy (anonymous):

I forgot what to do from here

OpenStudy (narissa):

k let me tag someone else thanks alot though

OpenStudy (narissa):

@skittles_for_life6422

OpenStudy (anonymous):

anytime

OpenStudy (narissa):

@dan815

OpenStudy (narissa):

@YoungStudier

OpenStudy (narissa):

dont u subtract 16 from - 28

OpenStudy (narissa):

@Data_LG2

OpenStudy (anonymous):